# Implicit Neural Representations for Robust Joint Sparse-View CT Reconstruction

### v1

@@ 5.1 Results
INR-Bayes $\textbf{30.13}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$ $\textbf{0.885}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $\textbf{31.20}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$ $\textbf{0.799}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{35.09}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.12}}$ $\textbf{0.602}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{28.24}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$ $\textbf{0.729}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$
@@
Applying to Unseen Data using Learned Prior. To evaluate the generalizability of the learned priors, we apply meta models or latent variables from the inter-object LungCT experiment to guide the reconstruction of new subjects in the dataset. Specifically, we select 5 consecutive slices from new objects, choosing slices from a similar location the prior has been trained. The prior information is solely utilized to guide the reconstruction and is not updated during the process. Table 3 shows that FedAvg fails to improve the reconstruction quality compared with singleINR, suggesting its learned meta neural representation struggles to generalize to unseen data. In contrast, both INR-Bayes and MAML effectively leverage their trained priors for improved reconstruction, with our method showing notably better metrics. Applying learned prior can also accelerate the training process. Figure 8 presents learning curves of different methods. All joint reconstruction methods converge faster than individual reconstruction. Initially, FedAvg converges the fastest, but as training progresses, both MAML and INR-Bayes surpass it. Additionally, the results reconfirm the robustness of INR-Bayes against overfitting, an issue that tends to challenge other compared methods.
@@
Table 3: Adaptation on new patients using priors learned from other patients compared to individual reconstruction.
@@
Broader Application. To assess the versatility of our proposed method, we expand its application to an unconventional domain by considering natural RGB images as 3-slice objects, akin to CT. Although a departure from traditional CT contexts, this experiment aims to test the methods’ generalizability.
The aggregate metrics for joint reconstruction across 10 groups, in total of 100 faces, are shown in Table 1. Notably, our approach surpasses other techniques by a considerable margin. This suggests not only its adaptability across diverse tasks but also underscores its potential applicability in other inverse problems, particularly where undersampling challenges reconstruction. A visual comparison in Figure 3 underscores our method’s prowess, revealing richer details and minimizing reconstruction artifacts. More results are relegated to Section D.2.
Computation. Our method takes 22.7 minutes on average for a $501\times 501$ reconstruction on A100 GPU, 15% longer than other INR methods. This computational overhead yields substantial gains in reconstruction quality and robustness across diverse settings. We also investigate the computational efficiency and performance trade-off in Section D.6.
@@ 6 Conclusions and Limitations
In this work, we study the possibility of improving the CT reconstruction quality through joint reconstruction and introduce a novel INR-based Bayesian framework. Through extensive experiments, our method has effectively showcased its ability to leverage the statistical regularities inherent in the sparse and noisy measurements of multiple objects to improve individual reconstructions. This capability allows our approach to outperform compared methods in reconstruction quality, robustness to overfitting as well as generalizability. While the primary focus of our method has been on joint CT reconstruction, its underlying principles hold potential applicability across a variety of inverse problems plagued by the challenges of sparse measurements.
@@ B.2 INRWild
We implement INRWild using the Siren network architecture (Sitzmann et al., 2020). Specifically, the static component, $\mathcal{G}_{\bm{\phi}}$, is represented using a standard 8-layer Siren network. In contrast, each of the transient parts, $\mathcal{H}_{\bm{w}_{j}}$, is characterized by a more compact 4-layer Siren network following the design from NeRFWild (Martin-Brualla et al., 2021). The optimization process ensures that the static and transient components are jointly optimized to ensure distinct representations, as articulated in Equation 3. A visual representation of the INRWild structure is provided in Figure 12.
@@ D.5 Joint Reconstruction across Temporal Phases in 4DCT
D.6 Computation Cost Analysis
@@
D.7 Comparison with Nerp
@@
D.8 Applying to Unseen Data using Different Learned Priors

### v2

@@ 5.1 Results
INR-Bayes $\textbf{30.13}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$ $\textbf{0.885}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $\textbf{31.20}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$ $\textbf{0.799}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{35.09}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.12}}$ $\textbf{0.869}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{28.24}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$ $0.729\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$
@@
Applying to Unseen Data using Learned Prior. To evaluate the generalizability of the learned priors, we apply meta models or latent variables from the inter-object LungCT experiment to guide the reconstruction of new subjects in the dataset. Specifically, we select 5 consecutive slices from new objects, choosing slices from a similar location (i.e. the similar axial slice index) the prior has been trained. The prior information is solely utilized to guide the reconstruction and is not updated during the process. Table 3 shows that FedAvg fails to improve the reconstruction quality compared with singleINR, suggesting its learned meta neural representation struggles to generalize to unseen data. In contrast, both INR-Bayes and MAML effectively leverage their trained priors for improved reconstruction, with our method showing notably better metrics. Applying learned prior can also accelerate the training process. Figure 8 presents learning curves of different methods. All joint reconstruction methods converge faster than individual reconstruction. Initially, FedAvg converges the fastest, but as training progresses, both MAML and INR-Bayes surpass it. Additionally, the results reconfirm the robustness of INR-Bayes against overfitting, an issue that tends to challenge other compared methods.
@@
Table 3: Adaptation on new patients using priors learned from other patients compared to individual reconstruction. The results are averaged by 10 new patients (5 slices each patient).
@@
Computation. Our method takes 22.7 minutes on average for a $501\times 501$ reconstruction on A100 GPU, 15% longer than other INR methods. This computational overhead yields noticeable gains in reconstruction quality and robustness across diverse settings. We also investigate the computational efficiency and performance trade-off in Section D.7.
@@ 6 Conclusions and Limitations
In this work, we study the possibility of improving the CT reconstruction quality through joint reconstruction and introduce a novel INR-based Bayesian framework. Through extensive experiments, our method has effectively showcased its ability to leverage the statistical regularities inherent in the sparse and noisy measurements of multiple objects to improve individual reconstructions. This capability allows our approach to outperform compared methods in reconstruction quality, robustness to overfitting as well as generalizability. While the primary focus of our method has been on joint CT reconstruction, its underlying principles hold potential applicability across a variety of inverse problems that are challenged by sparse measurements.
@@
Acknowledgments
Jiayang Shi, Daniël M. Pelt, K. Joost Batenburg and Matthew B. Blaschko are supported by the European Union H2020-MSCA-ITN-2020 under grant agreement no. 956172 (xCTing). Matthew B. Blaschko and Junyi Zhu received funding from the Flemish Government (AI Research Program) and the Research Foundation - Flanders (FWO) through project number G0G2921N.
@@ B.2 INRWild
We implement INRWild using the Siren network architecture (Sitzmann et al., 2020). Specifically, the static component, $\mathcal{G}_{\bm{\phi}}$, is represented using a standard 8-layer Siren network. In contrast, each of the transient parts, $\mathcal{H}_{\bm{w}_{j}}$, is characterized by a more compact 4-layer Siren network following the design from NeRFWild (Martin-Brualla et al., 2021). Each transient part is linked to a unique one-hot vector, which is embedded into a 16-dimensional transient feature $b_{j}$. This feature, alongside intermediate features $\mathcal{G}_{\bm{\phi}}^{\backslash r}(\bm{C})$ from the static network, serves as input for transient modeling. The transient feature $b_{j}$ is optimized during joint training.
The optimization process ensures that the static and transient components are jointly optimized to ensure distinct representations, as articulated in Equation 3. A visual representation of the INRWild structure is provided in Figure 12.
@@ B.4 Configuration of All Methods
Final Reconstructrion. After training, ours INR-Bayes yields a parameter distribution. Common methods for producing the final result based on this distribution include sampling the network parameters multiple times and averaging the network outputs, or using the mode of the distribution as the network parameters to produce the result (i.e., maximum a posteriori). In our work, we choose the latter approach, as it empirically yields better performance. For the baseline methods, we use the final trained model to produce the final reconstruction.
@@ D.5 Joint Reconstruction across Temporal Phases in 4DCT
D.6 Joint Reconstructions with Varied Object and Slice Configurations
To understand how different joint reconstruction configurations impact reconstruction quality, we examine four distinct setups: inter-object (10 objects, each contributing one slice), intra-object (one object contributing ten slices), and two intermediate configurations (two objects with five slices each, and five objects with two slices each). We conducted ten sets of experiments, with each set involving the joint reconstruction of the same ten slices across these configurations.
Table 5 presents the average PSNR and SSIM values obtained from these ten experimental groups. The results indicate that while the PSNR and SSIM values are generally consistent across different configurations, the intermediate configurations slightly outperform the others in terms of average metrics.
Figure 23 illustrates the learned priors from one exemplary set of reconstructions. Visually, the distinctions among the different configurations are minimal. The intra-object configuration (1×10) effectively captures the internal variations within a single object, aligning with our observation that a few intra-object slices are sufficient to model internal object changes. On the other hand, the 2×5 and 5×2 configurations, along with the inter-object (10×1) setup, depict priors that appear more distinct from the target slice, reflecting the influence of multiple objects in the joint reconstruction process. Notably, the 2×5 and 5×2 configurations exhibit slightly higher PSNR and SSIM values, suggesting that an optimal balance between similarity and diversity among jointly reconstructed objects can enhance the reconstruction quality of a specific target.
Config (${object}\times{slices}$) $1\times 10$ $2\times 5$ $5\times 2$ $10\times 1$
PSNR $35.56\stackrel{{\scriptstyle}}{{\scriptstyle\pm 1.03}}$ $35.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 1.08}}$ $35.66\stackrel{{\scriptstyle}}{{\scriptstyle\pm 1.07}}$ $35.59\stackrel{{\scriptstyle}}{{\scriptstyle\pm 1.02}}$
SSIM $0.876\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.025}}$ $0.877\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.025}}$ $0.878\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.025}}$ $0.876\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.025}}$
Table 5: Results from joint Reconstructions with varied object and slice configurations.
Figure 23: The reconstructions and learned priors with varied object and slice configurations.
D.7 Computation Cost Analysis
@@
D.8 Comparison with Nerp
@@
D.9 Applying to Unseen Data using Different Learned Priors
@@
D.10 Reconstruction Uncertainty
We employ a Bayesian framework to identify common patterns among jointly reconstructed objects. While the primary focus of this study is not on uncertainty quantification (Gal & Ghahramani, 2016; Angelopoulos et al., 2022), we demonstrate how INR-Bayes can facilitate this aspect within CT reconstructions. During training, each node approximates the posterior distribution of its weights as a Gaussian distribution. This approach allows us to sample the network parameters multiple times after training to quantify uncertainty. To illustrate this, we sample the network parameters ten times to generate ten distinct reconstructions in an intra-object setup using the LungCT dataset. We then compute the mean and variance of these reconstructions to show the expected reconstruction and characterize the associated uncertainty in Figure 26.
Figure 26: Mean and variance of 10 reconstructions by INR-Bayes.